# UAV traffic scene understanding: A cross-spectral guided approach and a unified benchmark

### 11 Mar 2026 v1

@@ 4.2 Data Construction Pipeline
Constructing a large-scale, cognitively rich, and robustly multimodal benchmark requires rigorous protocols to ensure both the visual fidelity of sensor data and the semantic integrity of linguistic annotations. To this end, we designed a systematic Human-in-the-Loop LLM-Empowered Construction Pipeline that minimizes manual annotation effort while expanding the structural diversity and cognitive complexity of the resulting queries. The construction workflow proceeds through three stages: Stage 1: Hardware-Synchronized Data Acquisition, Stage 2: Structured Attribute Annotation by Experts, and Stage 3: LLM-Empowered QA Generation.
@@
Table 2: Categorization of the 31 question types in the Traffic-VQA dataset, divided into cognitive, perceptual, and multimodal domains.
Question Type Object Type Specific Question Type Cognitive Traffic Participant Abnormal Traffic Condition, Vehicle Violations, Small Vehicle Violation, Pedestrian Violation, Vehicle Behavior, Small Vehicle Behavior, Deduce, Pedestrian Behavior, Traffic Participant Interaction, Road Condition Perceptual Traffic Participant Traffic Road, Vehicle, Pedestrian, Road Facility Total Object Compare, Presence, Location, Number, Shape, Most, Distribution, Residential, Agricultural, Industrial, Fog, Dark, Urban, Theme, UAV Modal Total Object Match, Modality
@@ 4.3 Dataset Analysis
Statistical Distribution of Question Types. The finalized dataset comprises 1,300,620 verified QA pairs derived from 8,180 image pairs. The queries are distributed across distinct functional categories. The most frequent types include Traffic Road (142,631 queries) and Modality (119,929 queries), reflecting the dataset’s focus on structural road understanding and cross-modal perception. Crucially, the dataset emphasizes higher-order cognition: Deduce alone includes 110,202 samples, and Location comprehension contributes 100,819 samples. Two other high-impact categories are Compare Vehicle (approximately 80k) and Vehicle Violation (42,674), both directly relevant to automated traffic surveillance. This distribution pushes evaluation beyond basic object presence toward situation understanding and rule-aware interpretation.
@@ 4.4 Dataset Challenges
In summary, Traffic-VQA provides a comprehensive platform for advancing multimodal cognitive intelligence in the domain of intelligent transportation systems.
@@ 5 Experiments
In this section, we present an empirical evaluation of the proposed CTCNet framework on the Traffic-VQA benchmark. We first describe the experimental configuration and implementation details (Section 5.1). We then provide a quantitative comparison against state-of-the-art (SOTA) MLLMs (Section 5.2), focusing on performance gains attributable to our cognitive and perceptual modules. Finally, ablation studies (Section 5.3) verify the individual contributions of the PGKE and QASC modules, complemented by a fine-grained analysis across question types.
@@ 5.1 Implementation Details
Model Configuration. We employ Qwen3-VL-8B-Instruct as the frozen backbone, selected for its strong visual-linguistic alignment capabilities. The input resolution for both optical and TIR images is set to $640\times 512$ pixels. To preserve the generalization capabilities of the pre-trained MLLM and ensure parameter efficiency, the core vision encoder and LLM backbone remain entirely frozen. Only the parameters of the Gated Parallel Residual Architecture—comprising the PGKE and QASC modules—are updated during training.
@@
Model Venue OA Traffic Cognition Traffic Perception Veh. Violation Veh. Behavior Veh. Ped. Violation Ped. Behavior Ped.
@@
CTCNet (Few-Shot) - 61.94 55.74 61.28 57.66 52.41 62.48 55.13 58.48 56.99
Qwen3-VL-8B (Finetuned) Ali 2025 79.598 80.546 75.296 76.747 75.65 79.766 76.496 71.245 69.661
Qwen3-VL-8B + CTCNet - 83.156 84.812 80.259 80.259 68.189 83.762 77.778 80.579 74.882
@@ 5.2 Quantitative Comparison
We benchmark the proposed CTCNet against two categories of baselines: (1) Zero-shot Open-source MLLMs (e.g., MiniCPM-V, GeoChat, Qwen3-VL Base); (2) Closed-source Commercial MLLMs (e.g., GPT-4o, Gemini-2.5-flash). The quantitative results, evaluated across single and multiple modalities, are summarized in Table 4.
@@ 5.2.1 Performance on Cognitive and Perceptual Tasks
As shown in Table 4, CTCNet achieves state-of-the-art performance across the board, with a consistent and substantial advantage. It is particularly strong in complex cognitive scenarios, effectively bridging the domain knowledge gap inherent in general-purpose models.
Superiority in Cognitive Understanding. The most critical evaluation domain is Traffic Cognitive Questions, which require an implicit understanding of traffic regulations combined with advanced contextual interpretation. While the Qwen3-VL-8B (Finetuned) baseline achieves a commendable cognitive accuracy of 80.55%—validating the utility of supervised domain adaptation—our fully integrated Qwen3-VL-8B + CTCNet architecture improves this to 84.81%, a margin of +4.26%. This improvement indicates that standard fine-tuning is insufficient to capture the multi-layered logic governing complex traffic violations. The advantage of CTCNet stems from the PGKE module, which retrieves expert prototypes from the external TRM, grounding the cognitive interpretation process in explicit, formalized domain knowledge rather than fragile statistical correlations.
Furthermore, CTCNet demonstrates strong performance in fine-grained violation detection, achieving notable accuracies in Vehicle Violation (80.26%) and Pedestrian Violation (77.78%). Notably, our method significantly outperforms leading commercial models such as GPT-4o (Cognitive: 75.28%) and Gemini-2.5-flash (Cognitive: 74.92%), demonstrating that a domain-specialized, knowledge-anchored module can substantially surpass large-scale general-purpose models in highly regulated vertical domains.
Robustness in Perceptual Tasks. For Traffic Perceptual Questions (e.g., object localization, presence detection), the fully fine-tuned CTCNet attains 80.26% accuracy, surpassing the corresponding fine-tuned baseline (75.30%) by nearly 5%. This improvement verifies that the QASC module effectively enhances feature distinctiveness, particularly for small and densely packed UAV targets. Even when constrained to the few-shot regime, our method (MUL OA: 61.94%) substantially outperforms all evaluated zero-shot open-source baselines (peak OA around 47.62%), validating that our lightweight residual injection strategy preserves general perception abilities while improving sensitivity to domain-specific semantic objects.
@@ 5.2.2 Evaluation of Multi-Spectral Robustness
In contrast, CTCNet consistently achieves its best performance in the MUL setting. Under the few-shot protocol, CTCNet improves the OPT baseline (47.56%) to a MUL OA of 61.94% (+14.38%). This result confirms that the QASC module successfully facilitates constructive interaction between modalities. By routing bidirectional attention toward the most informative regions of the TIR channel, the model effectively compensates for optical degradations (e.g., deep shadows, glare, or extreme low light) without incurring performance penalties from cross-modal feature conflict, ensuring reliable performance across all-weather UAV surveillance conditions.
@@ 5.3.1 Effectiveness of Proposed Modules
Table 5: Ablation study isolating the PGKE and QASC modules under the few-shot setting. Exp1 represents the unmodified baseline using simple feature concatenation. OA: Overall Accuracy, AA: Average Accuracy.
@@ 5.3.3 Fine-Grained Analysis by Question Type
CTCNet shows consistently strong performance across a diverse set of tasks, spanning both high-level cognitive analysis (e.g., "Abnormal Traffic Situation") and low-level perception queries (e.g., "Location"). These results suggest that the model generalizes well across task types and is a practical candidate for real-world UAV traffic surveillance scenarios.
@@ 6 Conclusion
In this paper, we construct a large-scale, unified benchmark dataset, Traffic-VQA, designed to advance all-weather UAV traffic cognitive understanding. Comprising 8,180 well-aligned OPT-TIR image pairs and over 1.3 million question-answer pairs, the dataset covers diverse environmental conditions and 31 distinct cognitive tasks. To address the limitations of existing methods regarding domain knowledge deficiency and cross-modality interference, we propose CTCNet. The CTCNet incorporates a PGKE module working in conjunction with an external TRM to inject domain-specific situational knowledge into visual features. Furthermore, the QASC module adaptively integrates complementary spectral information through dynamic, attention-driven context exchange. Extensive experiments on Traffic-VQA demonstrate the effectiveness of CTCNet, which significantly outperforms contemporary state-of-the-art MLLMs, particularly in high-level cognitive scenarios.
@@ Acknowledgement
This work was supported in part by the National Natural Science Foundation of China (No. 62306005, 62006002, and 62076003), in part by the Joint Funds of the National Natural Science Foundation of China (No. U20B2068), in part by the Natural Science Foundation of Anhui Province (No. 2208085J18 and 2208085QF192), and in part by the Natural Science Foundation of Anhui Higher Education Institution (No. 2022AH040014).

### 14 Mar 2026 v2

@@ 4.2 Data Construction Pipeline
Constructing a large-scale, cognitively rich, and robustly multi-modal benchmark requires rigorous protocols to ensure both the visual fidelity of sensor data and the semantic integrity of linguistic annotations. To this end, we designed a systematic Human-in-the-Loop LLM-Empowered Construction Pipeline that minimizes manual annotation effort while expanding the structural diversity and cognitive complexity of the resulting queries. The construction workflow proceeds through three stages: Stage 1: Hardware-Synchronized Data Acquisition, Stage 2: Structured Attribute Annotation by Experts, and Stage 3: LLM-Empowered QA Generation.
@@
Table 2: Categorization of the 31 question types in the Traffic-VQA dataset, divided into cognitive, perceptual, and multi-modal domains.
Question Type Object Type Specific Question Type Cognitive Traffic Participant Abnormal Traffic Situation, Vehicle Violations, Small Vehicle Violation, Pedestrian Violation, Vehicle Behavior, Small Vehicle Behavior, Deduce, Pedestrian Behavior, Traffic Participant Interaction, Road Condition Perceptual Traffic Participant Traffic Road, Vehicle, Pedestrian, Road Facility Total Object Compare, Presence, Location, Number, Shape, Most, Distribution, Residential, Agricultural, Industrial, Mist, Night, Urban, Theme, UAV Modal Total Object Match, Modality
@@ 4.3 Dataset Analysis
Statistical Distribution of Question Types. The finalized dataset comprises 1,301,466 verified QA pairs derived from 8,180 image pairs. The queries are distributed across distinct functional categories. The most frequent types include Traffic Road (142,631 queries) and Modality (119,929 queries), reflecting the dataset’s focus on structural road understanding and cross-modal perception. Crucially, the dataset emphasizes higher-order cognition: Deduce alone includes 110,202 samples, and Location comprehension contributes 100,819 samples. Two other high-impact categories are Compare Vehicle (approximately 80k) and Vehicle Violation (42,674), both directly relevant to automated traffic surveillance. This distribution pushes evaluation beyond basic object presence toward situation understanding and rule-aware interpretation.
@@ 4.4 Dataset Challenges
In summary, Traffic-VQA provides a comprehensive platform for advancing multi-modal cognitive intelligence in the domain of intelligent transportation systems.
@@ 5 Experiments
In this section, we present an empirical evaluation of the proposed MTCNet framework on the Traffic-VQA benchmark. We first describe the experimental configuration and implementation details (Section 5.1). We then provide a quantitative comparison against state-of-the-art (SOTA) MLLMs (Section 5.2), focusing on performance gains attributable to our cognitive and perceptual modules. Finally, ablation studies (Section 5.3) verify the individual contributions of the PGKE and QASC modules, complemented by a fine-grained analysis across question types.
@@ 5.1 Implementation Details
Model Configuration. We employ Qwen3-VL-8B as the frozen backbone, selected for its strong visual-linguistic alignment capabilities. The input resolution for both optical and TIR images is set to $640\times 512$ pixels. To preserve the generalization capabilities of the pre-trained MLLM and ensure parameter efficiency, the core vision encoder and LLM backbone remain entirely frozen. Only the parameters of the Gated Parallel Residual Architecture—comprising the PGKE and QASC modules—are updated during training.
@@
Model Source OA Traffic Cognition Traffic Perception Veh. Violation Veh. Behavior Veh. Ped. Violation Ped. Behavior Ped.
@@
MTCNet (Few-Shot) - 61.94 55.74 61.28 57.66 52.41 62.48 55.13 58.48 56.99
Qwen3-VL-8B (Finetuned) Ali 2025 79.6 80.55 75.3 76.75 75.65 79.77 76.5 71.25 69.66
Qwen3-VL-8B + MTCNet - 83.16 84.81 80.58 80.26 68.19 83.76 77.78 80.58 74.88
@@ 5.2 Quantitative Comparison
We benchmark the proposed MTCNet against two categories of baselines: (1) Zero-shot Open-source MLLMs (e.g., MiniCPM-V, GeoChat, Qwen3-VL Base); (2) Closed-source Commercial MLLMs (e.g., GPT-4o, Gemini-2.5-flash). The quantitative results, evaluated across single and multiple modalities, are summarized in Table 4.
@@ 5.2.1 Performance on Cognitive and Perceptual Tasks
As shown in Table 4, MTCNet achieves state-of-the-art performance across the board, with a consistent and substantial advantage. It is particularly strong in complex cognitive scenarios, effectively bridging the domain knowledge gap inherent in general-purpose models.
Superiority in Cognitive Understanding. The most critical evaluation domain is Traffic Cognitive Questions, which require an implicit understanding of traffic regulations combined with advanced contextual interpretation. While the Qwen3-VL-8B (Finetuned) baseline achieves a commendable cognitive accuracy of 80.55%—validating the utility of supervised domain adaptation—our fully integrated Qwen3-VL-8B + MTCNet architecture improves this to 84.81%, a margin of +4.26%. This improvement indicates that standard fine-tuning is insufficient to capture the multi-layered logic governing complex traffic violations. The advantage of MTCNet stems from the PGKE module, which retrieves expert prototypes from the external TRM, grounding the cognitive interpretation process in explicit, formalized domain knowledge rather than fragile statistical correlations.
Furthermore, MTCNet demonstrates strong performance in fine-grained violation detection, achieving notable accuracies in Vehicle Violation (80.26%) and Pedestrian Violation (77.78%). Notably, our method significantly outperforms leading commercial models such as GPT-4o (Cognitive: 75.28%) and Gemini-2.5-flash (Cognitive: 74.92%), demonstrating that a domain-specialized, knowledge-anchored module can substantially surpass large-scale general-purpose models in highly regulated vertical domains.
Robustness in Perceptual Tasks. For Traffic Perceptual Questions (e.g., object localization, presence detection), the fully fine-tuned MTCNet attains 80.26% accuracy, surpassing the corresponding fine-tuned baseline (75.30%) by nearly 5%. This improvement verifies that the QASC module effectively enhances feature distinctiveness, particularly for small and densely packed UAV targets. Even when constrained to the few-shot regime, our method (MUL OA: 61.94%) substantially outperforms all evaluated zero-shot open-source baselines (peak OA around 47.62%), validating that our lightweight residual injection strategy preserves general perception abilities while improving sensitivity to domain-specific semantic objects.
@@ 5.2.2 Evaluation of Multi-Spectral Robustness
In contrast, MTCNet consistently achieves its best performance in the MUL setting. Under the few-shot protocol, MTCNet improves the OPT baseline (47.56%) to a MUL OA of 61.94% (+14.38%). This result confirms that the QASC module successfully facilitates constructive interaction between modalities. By routing bidirectional attention toward the most informative regions of the TIR channel, the model effectively compensates for optical degradations (e.g., deep shadows, glare, or extreme low light) without incurring performance penalties from cross-modal feature conflict, ensuring reliable performance across all-weather UAV surveillance conditions.
@@ 5.3.1 Effectiveness of Proposed Modules
Table 5: Ablation study isolating the PGKE and QASC modules under the few-shot setting. Example 1 represents the unmodified baseline, where features are aggregated element-wise. OA: Overall Accuracy, AA: Average Accuracy.
@@ 5.3.3 Fine-Grained Analysis by Question Type
MTCNet shows consistently strong performance across a diverse set of tasks, spanning both high-level cognitive analysis (e.g., "Abnormal Traffic Situation") and low-level perception queries (e.g., "Location"). These results suggest that the model generalizes well across task types and is a practical candidate for real-world UAV traffic surveillance scenarios.
@@ 6 Conclusion
In this paper, we construct a large-scale, unified benchmark dataset, Traffic-VQA, designed to advance all-weather UAV traffic cognitive understanding. Comprising 8,180 well-aligned OPT-TIR image pairs and over 1.3 million question-answer pairs, the dataset covers diverse environmental conditions and 31 distinct cognitive tasks. To address the limitations of existing methods regarding domain knowledge deficiency and cross-modality interference, we propose MTCNet. The MTCNet incorporates a PGKE module working in conjunction with an external TRM to inject domain-specific situational knowledge into visual features. Furthermore, the QASC module adaptively integrates complementary spectral information through dynamic, attention-driven context exchange. Extensive experiments on Traffic-VQA demonstrate the effectiveness of MTCNet, which significantly outperforms contemporary state-of-the-art MLLMs, particularly in high-level cognitive scenarios.
@@ Acknowledgement
This work was supported in part by the National Natural Science Foundation of China (No. 62306005, 62006002, and 62076003), in part by the Natural Science Foundation of Anhui Province (No. 2208085J18 and 2208085QF192), and in part by the Natural Science Foundation of Anhui Higher Education Institution (No. 2022AH040014).